Need help with req B asapEMD Corporation manufactures two products, Product S and Product W. Product W is of fairly recent origin, having been developed as an attempt to enter a market closely related to that of Product W. Product W is the more complex of the two products, requiring 3 hours of direct labor time per unit to manufacture compared to 2 hour of direct labor time for Product S. Product W is produced on an automated production line Overhead is currently assigned to the products on the basis of direct-labor-hours. The company estimated it would incur $1,140,940 in manufacturing overhead costs and produce 10,500 units of Product W and 42,000 units of Product S during the current year. Unit cost for materials and direct labor are: Product SProduet w Direct material Direct labor 34 14 s 17 Required: a-1. Compute the predetermined overhead rate under the current method of allocation. a-2. Determine the unit product cost of each product for the current year. b. The companys overhead costs can be attributed to four major activities. These activities and the amount of overhead cost attributable to each for the current year are given below: Total Activit Total 3,770 693 18,690 1,615 Activity Cost Pool Total Cost 652,210 51,975 112,140 324,615 Products Product W Machine aetups required Purchase orders issued Machine-hours required Maintenance requests issued 1,510 531 7,900 737 2,260 162 10,790 878 $1,140,940 Using the data above and an activity-based costing approach, determine the unit product cost of each product for the current year.
Answer #1

B)

Step 1 :Calculation of overhead allocation to product and overhead cost per unit:

Activity Formula Product S Product W
Activity cost / Activity driver = Activity rate Activity used Overhead cost allocated (Activity used*activity rate) Activity used Overhead cost allocated (Activity used*activity rate)
Machine setup 652210 3770 173 1510 1510*173=261230 2260 2260*173=390980
Purchase order 51975 693 75 531 531*75=39825 162 162*75=12150
Machine hours required 112140 18690 6 7900 7900*6=47400 10790 10790*6=64740
Maintenance request issued 324615 1615 201 737 737*201=148137 878 878*201=176478
Total overhead cost allocated 496592 644348
Number of units 42000 10500
overhead cost per unit 496592/42000 =$ 11.82 644348/10500= 61.37

Calculation of unit cost:

Product S Product W
Direct material 17 34
Direct labor 8 14
overhead cost 11.82 61.37
Unit cost 36.82 109.37